OpenStudy (anonymous):

xy+x=12, dx/dt=-3, x=2, y=5 find dy/dt

OpenStudy (anonymous):

\[y+xy'+1=0\] put \[x=2,y=5\] solve for \(y'\)

OpenStudy (anonymous):

\[y \frac{ dx }{ dt }+x \frac{ dy }{ dt }+\frac{ dx }{ dt }=0\]

OpenStudy (anonymous):

and plug everything in (:

OpenStudy (anonymous):

ohhh waitt i think that's wrong

OpenStudy (anonymous):

Nevermind, it's correct

OpenStudy (anonymous):

what is dy/dt?

OpenStudy (anonymous):

by implicit differentiation and the chain rule

OpenStudy (anonymous):

\[(xy \prime+yx \prime)+x \prime=0\]

OpenStudy (anonymous):

\[y \prime \] is dy/dt and\[x \prime \] is dx/dt

OpenStudy (anonymous):

thanks so much ill would the answer be -3? or -9? those are two i have came up with

OpenStudy (anonymous):

It should be 9, not -9